A 65 g block of metal absorbs 5850 J of heat energy and the temperature of the block rises from 10 degrees Celsius to 41 degrees Celsius. What is the specific heat capacity of the metal?


1
Expert's answer
2021-08-10T05:03:34-0400

Q224396


A 65 g block of metal absorbs 5850 J of heat energy and the temperature of the block rise from 10 degrees Celsius to 41 degrees Celsius. What is the specific heat capacity of the metal?


Solution :


We are given the mass of the metal block, the block's initial temperature, and the block's final temperature after absorbing 5850 Joules of heat energy.


mass of metal block, m = 65 g


Initial temperature, Ti = 10 oC


Final temperature, Tf = 41 oC


Heat absorbed by the block, Q = 5850 Joules.


specific heat capacity of the metal, s = unknown



"Q = m \ s \ \Delta T"


"\Delta T = T_f - T_i = 41^o C - 10 ^o C = 31 ^o C"


plug all the information in the formula and solve for 's'.



"5850 \ J = 65 g \ * s * \ 31 ^o C"


"5850 \ J = 2015 \ g \ ^o C \ * \ s"

dividing both the side by 2015 g oC, we have



"\frac{5850 \ J}{2015 \ g \ ^o C } = \frac{\cancel{2015 \ g \ ^o C } \ * \ s }{\cancel{2015 \ g \ ^o C }}"

"\frac{5850 \ J}{2015 \ g \ ^o C } = \ s"

"2.903 \ J /g ^o C = \ s"


In the correct significant figure the answer is 2.9 J/g oC.


Hence the specific heat capacity of the metal is 2.9 J/g oC.